Find gof if f(x) =2x-1 and g(x) =x-2
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
Find gof if f(x) =2x-1 and g(x) =x-2

SOLUTION:

You want a rule for the composite function g(f(x)).

g(x) = x - 2

Now substitute 2x-1 for x in the function g(x).

g(f(x)) = (2x-1) - 2

Simplify.
g(f(x)) = 2x - 3

That's all there is to it! Here is a link to a nice explanation with diagrams and examples:
http://www.mathsisfun.com/sets/functions-composition.html

Good luck!
Mrs. Figgy